A shopper buys 3 oranges and 5 lemons for $10.26, while a second shopper buys 4 lemons and 6 oranges for $11.16. What is the price of each fruit?
I got 3x+5y=10.26 and 4x+6y=11.16
I gave the answer of x=2.88 y=.324 but i was wrong.

Answer by bucky(2189)

Here's a hint for future problems: use a letter for the variables that makes sense. In this problem instead of using x and y (easily confused) use O for oranges (that's letter O, not zero) and use L for lemons. Now write your equations.
.
For the three oranges and 5 lemons you get:
.
3*O + 5*L = 10.26
.
So far, no problem.
.
For the 4 lemons and 6 oranges you get:
.
6*O + 4*L = 11.16
.
Notice the difference in the second equation. You got the lemons and oranges reversed in order when you wrote 4x + 6y = 11.16.
.
Now you have the two equations:
.
3*O + 5*L = 10.26
6*O + 4*L = 11.16
.
Multiply the top equation (both sides and all terms) by 2 to get:
.
6*O + 10*L = 20.52 <--- First equation doubled
6*0 + 4*L = 11.16 <--- Second equation unchanged
.
Now subtract the two equations vertically and the answer becomes:
.
6*L = 9.36
.
Solve for the price of lemons by dividing by 6 to get:
.
L = 1.56
.
Now let's return to the second equation and substitute 1.56 for L. This equation then becomes:
.
6*O + (4*1.56) = 11.16
.
Multiply the terms in the parentheses and the equation becomes:
.
6*O + 6.24 = 11.16
.
Subtract 6.24 from both sides and you have:
.
6*O = 4.92
.
Divide both sides by 6 to get:
.
O = 0.82
.
So oranges are 82 cents each and lemons are $1.56 each.
.
Hope this helped you out a little bit. You obviously know what you are doing, but just made a simple mistake.